OpenStudy (anonymous):

A family of 3 children and 2 adults visited a health club. The club charges $y an hour for each child and $x an hour for each adult. If the family does not want to spend more than $30 an hour at the health club, create a graph that best models the situation.

OpenStudy (turingtest):

well we have two quantities, the the total amount spent, and the maximum they can spend the total amount spent is going to be totalCost = costOfKids + costOfAdults and we want the total cost less than $30, so totalCost <= $30 can you write this as an equation?

OpenStudy (turingtest):

*involving x and y I mean

OpenStudy (anonymous):

would it be $30=x+y

OpenStudy (turingtest):

not quite x and y are the *hourly* rates so we need another variable, t, to represent the time passed then xt=costOfKids make sense?

OpenStudy (anonymous):

i think so

OpenStudy (turingtest):

x is in dollars per hour, so dollars per hour times number of hours is how many dollars

OpenStudy (turingtest):

so what is an expression for the cost of the parents?

OpenStudy (anonymous):

would it be $30=xy

OpenStudy (turingtest):

no, that would just be multiplying the two rates together if your boss pays you $5/hour, and pays your friend $6/hour, do you make $30 ? I don't think so how much you make depends on the time you work

OpenStudy (turingtest):

the cost for an hourly service is (price per hour) x (number of hours) = total price

OpenStudy (turingtest):

for the kids, it will be (rate for kids) x (time) = cost of kids or xt for the adults it will be (rate for adults) x (time) = cost of adults how can you express this in symbols?

OpenStudy (anonymous):

3y=xt 2x=xt ?? idk im so stupid with math lol

OpenStudy (turingtest):

no, x is the rate for the kids im just asking about the cost for the parents (rate for parents) x (time) = cost for parents what is the rate for the parents?

OpenStudy (anonymous):

wait whats the time?

OpenStudy (turingtest):

t

OpenStudy (anonymous):

2x=t ?

OpenStudy (turingtest):

where did you get the 2 from?

OpenStudy (anonymous):

theres 2 adults

OpenStudy (turingtest):

let's just think about each adult first what is the cost of *one* adult?

OpenStudy (anonymous):

idkkkkk

OpenStudy (turingtest):

it is (rate for adults) x time what is the rate for adults?

OpenStudy (anonymous):

it doesnt say!!

OpenStudy (turingtest):

it does say - y

OpenStudy (turingtest):

$y an hour is the rate for adults

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

omg i dont know what you mean tho !

OpenStudy (turingtest):

cost of adult = (rate for adult) x time rate for adult = y time = t

OpenStudy (turingtest):

so cost of adult = ???? (in terms of y and t)

OpenStudy (anonymous):

y * t=cost of adults

OpenStudy (turingtest):

good, that is the cost of one adult how much then, is the cost of two adults?

OpenStudy (anonymous):

2y = cost of two adults ?

OpenStudy (turingtest):

great! how about the cost of the kids now?

OpenStudy (turingtest):

oh wait you forgot the time

OpenStudy (turingtest):

what is the cost of two adults again?

OpenStudy (anonymous):

2y * t =cost of two adults

OpenStudy (turingtest):

ok cool now what about the cost of the kids?

OpenStudy (anonymous):

x * t= cost of kids ?

OpenStudy (turingtest):

that is the cost of one kid how many kids are there?

OpenStudy (anonymous):

3x * t= cost

OpenStudy (turingtest):

great, so the total cost of kids and adults is what?

OpenStudy (anonymous):

2y * 3x * t = $30 ?

OpenStudy (turingtest):

to find the total of two things you add them, not multiply

OpenStudy (anonymous):

2y +3x+t=$30 ?

OpenStudy (turingtest):

well it turns out that $30 is the maximum *rate* they want to spend, so we don't need to consider the time anymore so you don't need t

OpenStudy (anonymous):

oh ok so just 2y+3x=$30

OpenStudy (turingtest):

almost, except it says they want to spend *no more than* $30 per hour so it shouldn't be =, but rather...?

OpenStudy (anonymous):

ok 2y + 3y < 30

OpenStudy (turingtest):

close,\[2x+3y\le30\]if it said that they wanted to spend *less than* 30 an hour you would use <, but they said *no more* than 30 an hour

OpenStudy (turingtest):

oh i got x and y mixed up

OpenStudy (turingtest):

3x+2y<=30

OpenStudy (anonymous):

yeah thats what i meant

OpenStudy (anonymous):

thamk youu!!